# S4ECG: Exploring the impact of long-range interactions for arrhythmia prediction

### v1

@@ 3 Results
We present a comprehensive evaluation of the S4ECG multi-epoch model across both in-distribution (ID) and out-of-distribution (OOD) scenarios. Our results demonstrate consistent and substantial improvements of multi-epoch models over conventional single-epoch approaches across all evaluated datasets and metrics, with statistical significance confirmed using a patient-level paired bootstrap (10,000 resamples; 95% confidence intervals, CIs).
@@
3.1 Influence of the number of input epochs
Systematic evaluation of multi-epoch model performance reveals clear optimal ranges for the number of input epochs, with distinct patterns emerging across different training datasets and evaluation scenarios. Figure 2 illustrates the performance trends across different temporal contexts.
ID performance trends. For models trained on Icentia11k (Table 3), the single-epoch baseline achieves a macro-AUROC of 0.970, establishing a strong foundation. The S4-based single-epoch model already outperforms the ResNet baseline by 0.4% (0.966 vs. 0.970), confirming S4’s architectural advantages before considering multi-epoch benefits in line with previous investigations [15]. Multi-epoch models demonstrate consistent improvements, with optimal performance achieved at 30 input epochs (macro-AUROC: 0.980, +1.0% improvement). Uncertainty estimates are on the order of $10^{-4}$ level across models. This represents improvements across all rhythm classes. At a fixed AF sensitivity of 0.9, specificity improves from 0.903 (single-epoch S4) to 0.987 (30 epochs).
For LTAFDB training (Table 4), the multi-epoch advantage is even more pronounced. The single-epoch model achieves a modest macro-AUROC of 0.903, representing an 4.7% improvement over the ResNet baseline (0.862), which again confirms S4’s superiority for this challenging dataset before any multi-epoch modeling. Multi-epoch models show dramatic improvements from as few as two input epochs. Peak performance occurs at 10 input epochs (macro-AUROC: 0.968, +7.3% improvement), with exceptional gains in atrial fibrillation detection (AF: 0.832 to 0.984, +18.3%). Notably, the performance remains consistently high across the 10-40 epoch range (macro-AUROC: 0.961-0.968), with different rhythm classes achieving their optimal performance at different points within this stable range: normal rhythm (N) at 20 epochs (0.995), and SVTA at 40 epochs (0.929). Correspondingly, AF specificity rises from 0.979 (single-epoch S4) to 0.998 (30 epochs). Among these models, the 20-epoch configuration achieves statistically equivalent performance to the best model (macro-AUROC: 0.967), demonstrating the robustness of the multi-epoch approach across this temporal range and providing flexibility in clinical deployment scenarios.
@@ 3.1 Influence of the number of input epochs
Model Type Input Epochs AUROC Spec.
@@
Table 4: LTAFDB in-distribution performance (model trained and evaluated on LTAFDB): Confirming S4’s superiority over CNN (ResNet) established in Table 3, and combined with prior evidence of S4’s advantages over LSTM and Transformers [13], justifying our focus on S4-based architectures. Underlined bold macro-AUROC: best performing model. Bold macro-AUROC: statistically equivalent to best model. Bold class-AUROC: highest values within each class. Spec.: AF specificity at sensitivity of 0.9.
Model Type Input Epochs AUROC Spec.
@@
Single-epoch
Model(Resnet)
[…]
Single-epoch
Model(S4)
[…]
Multi-epoch
Model(S4ECG)
@@
Class-specific analysis The rhythm-specific improvements reveal important insights into the clinical value of temporal context. AF detection shows the most consistent improvements, with multi-epoch models often reaching AUROC $\geq 0.98$ in ID settings and $\geq 0.95$ in OOD evaluations. This finding aligns with the clinical understanding that atrial fibrillation episodes often exhibit characteristic temporal patterns that extend beyond individual 30-second epochs. In contrast, normal rhythm classification shows more modest but consistent improvements, suggesting that the temporal context helps distinguish true normal rhythms from transient artifacts or brief arrhythmic episodes.
Moderate sequence length advantages As shown in Figure 2(a), the performance exhibits a characteristic inverted U-shape, with diminishing returns observed beyond 40 epochs, suggesting that moderate input sequence lengths are optimal for capturing temporal dependencies without overfitting, a finding that aligns with theoretical insights from sleep staging research demonstrating that moderate-length temporal windows provide the best balance between context richness and model generalization [24]. The consistent optimal performance in the 20-40 epoch range reflects this principle, where sufficient temporal context is provided to capture arrhythmic patterns without introducing excessive noise or computational complexity. This finding has important implications for practical deployment, as it suggests that effective arrhythmia detection does not require excessively long monitoring windows, making the approach suitable for real-time clinical applications. A qualitatively similar effect was observed in a recent study [25] upon studying the optimal input size for interpreting 10-second 12-lead ECGs, which suggested that these signals do not carry long-range interactions beyond 2.5-3 seconds. While the latter can be assumed to be stationary across 10 seconds, the results achieved in this work suggest that diagnostically relevant long-range interactions for arrhythmia detection in non-stationary long-term ECGs remain limited to time frames around 10-15 minutes.
@@ 3.2 OOD evaluation
Dataset Model Type Input Epochs AUROC Spec.
@@
AFDB
Single-epoch
Model
[…]
Multi-epoch
Model(S4ECG)
@@
MITDB
Single-epoch
Model
[…]
Multi-epoch
Model(S4ECG)
@@
LTAFDB
Single-epoch
Model
[…]
Multi-epoch
Model(S4ECG)
@@
The OOD evaluation on three external datasets provides crucial insights into model generalization capabilities and reveals that multi-epoch models demonstrate superior robustness across diverse clinical scenarios and acquisition protocols.
Cross-dataset generalization from Icentia11k Models trained on the large-scale Icentia11k dataset exhibit remarkable robustness when evaluated on external datasets (AFDB, MITDB, and LTAFDB) as shown in Table 5. For the AFDB evaluation, single-epoch models achieve a macro-AUROC of 0.8718, while multi-epoch models reach peak performance at 30 input epochs (macro-AUROC: 0.9328, +7.0% improvement). Correspondingly, AF specificity at sensitivity 0.9 improves strongly from 0.9572 (single-epoch) to near-perfect 0.9998 (30 epochs), demonstrating exceptional reduction in false positives. The improvement is particularly striking for atrial flutter detection (AFLT: 0.6899 to 0.8190, +18.7%), demonstrating the value of temporal context for detecting this challenging arrhythmia type in OOD settings.
Similarly, on MITDB evaluation, multi-epoch models show consistent improvements, with optimal performance again at 30 input epochs (macro-AUROC: 0.9401 vs. 0.8426 for single-epoch, +11.6% improvement) and the 40-epoch configuration achieves statistically equivalent performance (macro-AUROC: 0.9382), demonstrating robustness across this temporal range. The AF detection specificity shows substantial improvement from 0.7962 (single-epoch) to 0.9743 (40 epochs), reflecting enhanced clinical utility through reduced false alarms. The substantial improvement in normal rhythm classification (N: 0.6983 to 0.8465, +21.2%) suggests that multi-epoch models are particularly effective at maintaining specificity in challenging OOD scenarios where signal characteristics may differ significantly from training data.
Cross-dataset evaluation from Icentia11k training to LTAFDB testing reveals exceptional generalization capabilities, with the 30-epoch model achieving a macro-AUROC of 0.9767 (+5.5% over single-epoch). Most notably, AF specificity improves from 0.718 (single-epoch) to 0.9884 (20 epochs), representing a dramatic 37.6% increase that substantially reduces false positive burden. AF detection shows improvement from the single-epoch baseline (AF: 0.9427 to 0.9748, +3.4%), demonstrating that the temporal patterns learned from Icentia11k’s diverse population effectively generalize to detect atrial fibrillation in LTAFDB’s AF-focused dataset. The consistently high performance across the multi-epoch range mirrors the stability patterns observed in ID evaluation, suggesting that the optimal temporal window characteristics are robust across different dataset domains and patient populations.
@@ 3.3 Quantitative performance analysis
Performance stability across epoch counts The multi-epoch advantage is consistently substantial across all evaluation scenarios. In-distribution improvements range from 1.0% (Icentia11k) to 7.3% (LTAFDB) in macro-AUROC, while OOD improvements are even more pronounced, ranging from 5.5% to 11.6%. These improvements represent clinically meaningful enhancements in diagnostic accuracy, particularly for rare but critical arrhythmias. Analysis of performance across different epoch counts reveals remarkable stability in the 20-40 epoch range, with peak performance consistently achieved at 30 epochs for Icentia11k-trained model ID and OOD validation. This stability suggests robust optimal hyperparameter selection and practical deployment considerations, as moderate variations in sequence length do not dramatically impact performance. The observed performance curve, characterized by rapid improvements from single-epoch to moderate multi-epoch models followed by gradual degradation at excessive sequence lengths, corroborates the theoretical framework proposed by Wang and Strodthoff [24] regarding the optimal temporal window for physiological signal analysis. This pattern reflects the fundamental trade-off between capturing meaningful long-range dependencies and avoiding the curse of dimensionality in sequence modeling.
Rhythm-specific insights Atrial fibrillation detection shows the most consistent improvements across all scenarios, with multi-epoch models often achieving AUROC $\geq 0.98$ in ID settings and $\geq 0.95$ in OOD evaluations. This finding supports the clinical intuition that atrial fibrillation episodes exhibit characteristic temporal signatures that extend beyond individual epochs. Normal rhythm classification improvements, while more modest (typically 1-3%), are consistently present and particularly valuable for maintaining specificity in clinical applications.
Cross-domain validation of moderate sequence lengths Our results provide strong empirical validation of the theoretical insights from the sleep staging domain [24], demonstrating that the principle of moderate sequence length optimization generalizes across different physiological monitoring applications. The consistent 20-40 epoch optimal range observed in our ECG arrhythmia detection task mirrors the findings in sleep stage classification, where similar moderate temporal windows proved most effective. This cross-domain consistency suggests a fundamental characteristic of physiological time series analysis, where intermediate-length sequences provide the optimal balance between temporal context richness and model generalization capability. Such findings have broader implications for the design of temporal models in biomedical signal processing, supporting the adoption of moderate sequence lengths as a general principle rather than a task-specific optimization.
@@ 3.4 Qualitative insights and clinical impact
Figure 3: Qualitative comparison of model predictions on a continuous atrial fibrillation episode from LTAFDB. Top to bottom: ground truth annotation, multi-epoch model (30-epoch input), single-epoch model, and enlarged detail. The multi-epoch model maintains temporal coherence matching the ground truth, while the single-epoch model produces fragmented predictions with spurious interruptions, demonstrating the superior temporal consistency of the multi-epoch approach for arrhythmia burden estimation.
To provide deeper insights into multi-epoch model behavior and temporal pattern recognition capabilities, we present qualitative analysis of individual ECG recordings showing how the S4ECG model, which is trained on Icentia11k at 30 epochs input lengths, processes extended temporal sequences. Figure 3 presents a representative example from a test recording in the LTAFDB dataset, illustrating the multi-epoch model’s superior temporal coherence and arrhythmia burden estimation capabilities.
The visualization includes analysis of AF burden, comparing actual and predicted AF loads across the extended monitoring period from this LTAFDB recording. The multi-epoch model demonstrates superior accuracy in estimating the overall proportion of time spent in AF, which is a clinically important metric for patient risk stratification and treatment decisions.
Unlike single-epoch models that may produce inconsistent predictions across adjacent time segments, the multi-epoch S4ECG model generates temporally coherent predictions that better align with the underlying physiological patterns. The extended temporal context enables the model to maintain consistency across rhythm transitions and reduces fragmented episode detection, as demonstrated in this LTAFDB test recording.
@@ 4.1 Technical implications
Summary This work presents, to our knowledge, the first systematic, cross-dataset investigation of multi-epoch deep learning approaches for ECG arrhythmia detection, introducing S4ECG, a novel architecture that leverages structured state space models to capture long-range temporal dependencies in cardiac rhythm analysis. Our systematic evaluation across four major ECG databases demonstrates consistent and substantial improvements of multi-epoch models over conventional single-epoch approaches, with particularly notable gains in out-of-distribution scenarios.
Multi-epoch paradigm The most significant finding is the consistent optimal performance achieved in the 20-40 epoch range (10-20 minutes of ECG data), suggesting fundamental characteristics of cardiac rhythm analysis that extend beyond dataset-specific artifacts. Multi-epoch models achieve statistically significant improvements ranging from 1.0% to 11.6% in macro-AUROC across different evaluation scenarios, with particularly striking gains for challenging arrhythmia types such as atrial flutter (+18.7% improvement in OOD settings). Similarly, multi-epoch models, in comparison to single-epoch models, show a substantially increased specificity at a fixed sensitivity level of 0.9, indicating a substantial reduction in false positive predictions. The stability of these improvements across diverse datasets, acquisition protocols, and patient populations provides strong evidence for the generalizability of multi-epoch approaches.
We envision that this work will contribute to a paradigm shift in arrhythmia detection algorithm design, encompassing both temporal modeling evolution—from single-epoch analysis toward temporally-aware multi-epoch approaches—and architectural advancement from traditional CNN/LSTM frameworks toward efficient structured state space models that better align with clinical practice and the inherent temporal nature of cardiac arrhythmias.

### v2

@@ 3 Results
We present a comprehensive evaluation of the S4ECG multi-window model across both ID and OOD scenarios. Our results demonstrate consistent and substantial improvements of multi-window models over conventional single-window approaches across all evaluated datasets and metrics, with statistical significance confirmed using a patient-level paired bootstrap (10,000 resamples; 95% confidence intervals, CIs).
@@
3.1 Influence of the number of input windows
Systematic evaluation of multi-window model performance reveals clear optimal ranges for the number of input windows, with distinct patterns emerging across different training datasets and evaluation scenarios. Figure 2 illustrates the performance trends across different temporal contexts.
ID performance trends. For models trained on Icentia11k (Table 3), the S4-based single-window baseline achieves a macro-AUROC of 0.970, establishing a high-performance baseline. The S4-based single-window model already outperforms the ResNet baseline by 0.4% (0.9702 vs. 0.9663), confirming S4’s architectural advantages before considering multi-window benefits in line with previous investigations (Al-Masud et al., 2025). Multi-window models demonstrate consistent improvements, with optimal performance achieved at 30 input windows (macro-AUROC: 0.9800, +1.0% improvement). Uncertainty estimates are on the order of $10^{-4}$ across models. This represents improvements across all rhythm classes. At a fixed AF sensitivity of 0.9, specificity improves from 0.9033 (single-window S4) to 0.9869 (30 windows).
For LTAFDB training (Table 4), the multi-window advantage is even more pronounced. The single-window model achieves a modest macro-AUROC of 0.9029, representing a 4.7% improvement over the ResNet baseline (0.8621), which again confirms S4’s superiority for this challenging dataset before any multi-window modeling. Multi-window models show substantial improvements from as few as two input windows. Peak performance occurs at 10 input windows (macro-AUROC: 0.9684, +7.3% improvement), with exceptional gains in atrial fibrillation detection (AF: 0.8319 to 0.9841, +18.3%). Notably, the performance remains consistently high across the 10-40 window range (macro-AUROC: 0.9612-0.9684), with different rhythm classes achieving their optimal performance at different points within this stable range: normal rhythm (N) at 20 windows (0.9950), and SVTA at 40 windows (0.9285). Correspondingly, AF specificity rises from 0.9794 (single-window S4) to 0.9983 (30 windows). Among these models, the 20-window configuration achieves performance that is statistically equivalent to the best model (macro-AUROC: 0.9664), demonstrating the robustness of the multi-window approach across this temporal range and providing flexibility in clinical deployment scenarios.
@@ 3.1 Influence of the number of input windows
Model Type Input Windows AUROC Spec.
@@
Table 4: LTAFDB in-distribution performance (model trained and evaluated on LTAFDB): Confirming S4’s superiority over CNN (ResNet) established in Table 3, and combined with prior evidence of S4’s advantages over LSTM and Transformers (Wang and Strodthoff, 2025a), justifying our focus on S4-based architectures. Underlined bold macro-AUROC: best performing model. Bold macro-AUROC: statistically equivalent to best model. Bold class-AUROC: highest values within each class. Spec.: AF specificity at sensitivity of 0.9.
Model Type Input Windows AUROC Spec.
@@
[…]
Class-specific analysis The rhythm-specific improvements reveal important insights into the clinical value of temporal context. AF detection shows the most consistent improvements, with multi-window models often reaching AUROC $\geq 0.98$ in ID settings and $\geq 0.95$ in OOD evaluations. This finding aligns with the clinical understanding that atrial fibrillation episodes often exhibit characteristic temporal patterns that extend beyond individual 30-second windows. In contrast, normal rhythm classification shows more modest but consistent improvements, suggesting that the temporal context helps distinguish true normal rhythms from transient artifacts or brief arrhythmic episodes.
Moderate sequence length advantages As illustrated in Figure 2(a), the performance trajectory is non-monotonic, characterized by an initial rapid improvement followed by a plateau. A slight performance degradation is observed beyond 40 windows. We hypothesize that this decline primarily reflects optimization challenges than architectural limitations: although S4 layers can theoretically capture long-range dependencies with linear complexity, longer sequences induce a more complex loss landscape with numerous local minima, making gradient-based optimization more difficult and sensitive to initialization. This interpretation is supported by prior work on sleep staging (Wang and Strodthoff, 2024), which demonstrated that curriculum training—progressively extending the sequence length during training—recovers the performance lost under direct long-sequence training, suggesting that the limitation lies in the optimization trajectory rather than in the model architecture or data characteristics. Secondary contributing factors may include diminishing diagnostic returns beyond the physiologically meaningful timescale of approximately 10–20 minutes for most arrhythmias. Taken together, these effects suggest that moderate input sequence lengths (20–40 windows) capture diagnostically relevant dependencies while avoiding the optimization difficulties associated with very long sequences. A qualitatively similar effect was observed in a recent study (Mehari and Strodthoff, 2023) upon studying the optimal input size for interpreting 10-second 12-lead ECGs, which suggested that these signals do not carry long-range interactions beyond 2.5-3 seconds. While the latter can be assumed to be stationary across 10 seconds, the results achieved in this work suggest that diagnostically relevant long-range interactions for arrhythmia detection in non-stationary long-term ECGs remain limited to time frames around 10-15 minutes.
@@ 3.2 OOD evaluation
Dataset Model Type Input Windows AUROC Spec.
@@
[…]
The OOD evaluation on three external datasets provides crucial insights into model generalization capabilities and reveals that multi-window models demonstrate superior robustness across diverse clinical scenarios and acquisition protocols.
Cross-dataset generalization from Icentia11k Models trained on the large-scale Icentia11k dataset exhibit remarkable robustness when evaluated on external datasets (AFDB, MITDB, and LTAFDB) as shown in Table 5. For the AFDB evaluation, single-window models achieve a macro-AUROC of 0.8718, while multi-window models reach peak performance at 30 input windows (macro-AUROC: 0.9328, +7.0% improvement). Correspondingly, AF specificity at sensitivity 0.9 improves strongly from 0.9572 (single-window) to 0.9998 (30 windows), demonstrating exceptional reduction in false positives. The improvement is particularly striking for atrial flutter detection (AFLT: 0.6899 to 0.8190, +18.7%), demonstrating the value of temporal context for detecting this challenging arrhythmia type in OOD settings.
Similarly, on MITDB evaluation, multi-window models show consistent improvements, with optimal performance again at 30 input windows (macro-AUROC: 0.9401 vs. 0.8426 for single-window, +11.6% improvement) and the 40-window configuration achieves statistically equivalent performance (macro-AUROC: 0.9382), demonstrating robustness across this temporal range. The AF detection specificity shows substantial improvement from 0.7962 (single-window) to 0.9743 (40 windows), reflecting enhanced clinical utility through reduced false alarms. The substantial improvement in normal rhythm classification (N: 0.6983 to 0.8465, +21.2%) suggests that multi-window models are particularly effective at maintaining specificity in challenging OOD scenarios where signal characteristics may differ significantly from training data.
Cross-dataset evaluation from Icentia11k training to LTAFDB testing reveals robust generalization capabilities, with the 30-window model achieving a macro-AUROC of 0.9767 (+5.5% over single-window). Most notably, AF specificity improves from 0.718 (single-window) to 0.9884 (20 windows), representing a dramatic 37.6% increase that substantially reduces false positive burden. AF detection shows improvement from the single-window baseline (AF: 0.9427 to 0.9748, +3.4%), demonstrating that the temporal patterns learned from Icentia11k’s diverse population effectively generalize to detect atrial fibrillation in LTAFDB’s AF-focused dataset. The consistently high performance across the multi-window range mirrors the stability patterns observed in the ID evaluation, suggesting that the optimal temporal window characteristics are robust across different dataset domains and patient populations.
@@ 3.3 Quantitative performance analysis
Performance stability across window counts The multi-window advantage is consistently substantial across all evaluation scenarios. In-distribution improvements range from 1.0% (Icentia11k) to 7.3% (LTAFDB) in macro-AUROC, while OOD improvements are even more pronounced, ranging from 5.5% to 11.6%. These improvements represent clinically meaningful enhancements in diagnostic accuracy, particularly for rare but critical arrhythmias. Analysis of performance across different window counts reveals remarkable stability in the 20-40 window range, with peak performance consistently achieved at 30 windows for Icentia11k-trained model ID and OOD validation. This stability suggests robust optimal hyperparameter selection and practical deployment considerations, as moderate variations in sequence length do not dramatically impact performance. The observed performance curve, characterized by rapid improvements from single-window to moderate multi-window models followed by gradual degradation at excessive sequence lengths, corroborates the theoretical framework proposed by Wang and Strodthoff (Wang and Strodthoff, 2024) regarding the optimal temporal window for physiological signal analysis. This pattern reflects the fundamental trade-off between capturing meaningful long-range dependencies and avoiding the curse of dimensionality in sequence modeling.
Rhythm-specific insights Atrial fibrillation detection shows the most consistent improvements across all scenarios, with multi-window models often achieving AUROC $\geq 0.98$ in ID settings and $\geq 0.95$ in OOD evaluations. This finding supports the clinical intuition that atrial fibrillation episodes exhibit characteristic temporal signatures that extend beyond individual windows. Normal rhythm classification improvements, while more modest (typically 1-3%), are consistently present and particularly valuable for maintaining specificity in clinical applications.
Cross-domain validation of moderate sequence lengths Our results provide strong empirical validation of the theoretical insights from the sleep staging domain (Wang and Strodthoff, 2024), demonstrating that the principle of moderate sequence length optimization generalizes across different physiological monitoring applications. The consistent 20-40 window optimal range observed in our ECG arrhythmia detection task mirrors the findings in sleep stage classification, where similar moderate temporal windows proved most effective. This cross-domain consistency suggests a fundamental characteristic of physiological time series analysis, where intermediate-length sequences provide the optimal balance between temporal context richness and model generalization capability. Such findings have broader implications for the design of temporal models in biomedical signal processing, supporting the adoption of moderate sequence lengths as a general principle rather than a task-specific optimization.
@@ 3.4 Qualitative insights and clinical impact
Figure 3: Qualitative comparison of model predictions on a continuous atrial fibrillation episode from LTAFDB. Top to bottom: ground truth annotation, multi-window model (30-window input), single-window model, and enlarged detail. The multi-window model maintains temporal coherence matching the ground truth, while the single-window model produces fragmented predictions with spurious interruptions, demonstrating the superior temporal consistency of the multi-window approach for arrhythmia burden estimation.
To provide deeper insights into multi-window model behavior and temporal pattern recognition capabilities, we present qualitative analysis of individual ECG recordings showing how the S4ECG model, which is trained on Icentia11k at 30 windows input length, processes extended temporal sequences. Figure 3 presents a representative example from a test recording in the LTAFDB dataset, illustrating the multi-window model’s superior temporal coherence and arrhythmia burden estimation capabilities.
The visualization includes analysis of AF burden, comparing actual and predicted AF loads across the extended monitoring period from this LTAFDB recording. The multi-window model demonstrates superior accuracy in estimating the overall proportion of time spent in AF, which is a clinically important metric for patient risk stratification and treatment decisions.
Unlike single-window models that may produce inconsistent predictions across adjacent time segments, the multi-window S4ECG model generates temporally coherent predictions that better align with the underlying physiological patterns. The extended temporal context enables the model to maintain consistency across rhythm transitions and reduces fragmented episode detection, as demonstrated in this LTAFDB test recording.
@@ 4.1 Technical implications
Summary This work presents, to our knowledge, the first systematic, cross-dataset investigation of multi-window deep learning approaches for ECG arrhythmia detection, introducing S4ECG, a novel architecture that leverages structured state space models to capture long-range temporal dependencies in cardiac rhythm analysis. Our systematic evaluation across four major ECG databases demonstrates consistent and substantial improvements of multi-window models over conventional single-window approaches, with particularly notable gains in out-of-distribution scenarios.
Multi-window paradigm The most significant finding is the consistent optimal performance achieved in the 20-40 window range (10-20 minutes of ECG data), suggesting fundamental characteristics of cardiac rhythm analysis that extend beyond dataset-specific artifacts. Multi-window models achieve statistically significant improvements ranging from 1.0% to 11.6% in macro-AUROC across different evaluation scenarios, with particularly striking gains for challenging arrhythmia types such as atrial flutter (+18.7% improvement in OOD settings). Similarly, multi-window models, in comparison to single-window models, show a substantially increased specificity at a fixed sensitivity level of 0.9, indicating a substantial reduction in false positive predictions. The stability of these improvements across diverse datasets, acquisition protocols, and patient populations provides strong evidence for the generalizability of multi-window approaches.
We envision that this work will contribute to a methodological transition in arrhythmia detection algorithm design, encompassing both temporal modeling evolution—from single-window analysis toward temporally-aware multi-window approaches—and architectural advancement from traditional CNN/LSTM frameworks toward efficient structured state space models that better align with clinical practice and the inherent temporal nature of cardiac arrhythmias.